# Convergence Rates of Constrained Expected Improvement

### v1

@@ Abstract
Constrained Bayesian optimization (CBO) methods have seen significant success in black-box optimization with constraints, and one of the most commonly used CBO methods is the constrained expected improvement (CEI) [1] algorithm. CEI is a natural extension of the expected improvement (EI)  [2] when constraints are incorporated. However, the theoretical convergence rate of CEI has not been established. In this work, we study the convergence rate of CEI by analyzing its simple regret upper bound. First, we show that when the objective function $f$ and constraint function $c$ are assumed to each lie in a reproducing kernel Hilbert space (RKHS), CEI achieves the convergence rates of $\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)\right)\ \text{and }\ \mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)\right)$ for the commonly used squared exponential and Matérn kernels, respectively. Second, we show that when $f$ and $c$ are assumed to be sampled from Gaussian processes (GPs), CEI achieves the same convergence rates with a high probability. Numerical experiments are performed to validate the theoretical analysis.
@@ 1 Introduction
Bayesian optimization (BO) is an efficient method for optimizing expensive black-box functions without derivatives. It leverages probabilistic surrogate models, most commonly Gaussian processes (GPs), to balance exploration and exploitation in the search for optimal solutions [3]. BO has found widespread success in diverse fields such as structural design [4], machine learning hyperparameter tuning [5], robotics [6], inertial confinment fusion design [7], process parameter optimization of additive manufacturing [8], etc.
@@
While traditional BO is typically applied to unconstrained settings, many real-world problems involve black-box constraints that must be satisfied. This has motivated growing interest in constrained Bayesian optimization (CBO), where surrogate models are also constructed for constraint functions [9] that are complex and expensive to evaluate, making CBO especially valuable in applications like engineering design [10] and automated machine learning  [11]. One of the very key difference between unconstrained and constrained optimization is that the feasible region for constrained optimization problem consists of the search space where all constraints must be satisfied. A general formulation of the constrained BO problem is:
$\underset{{\bm{x}}\in C}{\text{minimize}}$ $f({\bm{x}}),$ subject to $c({\bm{x}})\leq 0,$ (1)
@@
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is the objective function, and $c:\mathbb{R}^{d}\to\mathbb{R}^{m}$ are the constraint functions. Both are defined on a compact input space $C\subset\mathbb{R}^{d}$. The objective and the constraint functions are both expensive black-box functions, that can only be evaluated through expensive physical or computer experiments. Throughout this paper, we consider the noise-free setting for both the objective and the constraints, i.e., the function evaluations are deterministic and the true function values can be observed. In addition, a single constraint is considered, i.e., $m=1$, for simplicity of presentation. We note that our analysis can be easily extended to multiple constraints (see Remark 3.15 for details).
@@
Broadly, CBO methods can be categorized into implicit and explicit approaches [12]. Implicit methods modify standard acquisition functions to incorporate constraints via merit functions or feasibility weights. Explicit methods estimate the feasible region directly and restrict search to this region. Among these, the constrained expected improvement (CEI) [13, 14, 1] stands out as one of the most basic and widely adopted methods. CEI is a natural extension of the well-known expected improvement (EI) function [2], where the acquisition function is computed as the product of EI and the probability of feasibility. Thanks to this simple and interpretable formulation, CEI has been successfully applied across domains, and it remains one of the default choices in many constrained BO software packages  [15].
@@
Despite its empirical popularity, the theoretical understanding of CEI lags behind. In contrast, unconstrained EI has been more extensively studied. Under a frequentist assumption where the objective $f$ lies in a reproducing kernel Hilbert space (RKHS), [16] established the convergence rate of EI by deriving the simple regret upper bound. Other works explored the density of sampled sequences [17] or connections between EI and optimal computing budget allocation [18]. However, convergence rates (i.e., simple regret upper bound) for CEI have not been rigorously established—neither under frequentist nor under Bayesian settings. Here, Bayesian setting means the objective $f$ is a random function sampled from a GP.
@@
In this paper, we provide the first theoretical convergence rates for CEI, focusing on simple regret upper bounds under both the frequentist and Bayesian settings. Our convergence rates provide practitioners theoretical assurance for the practical deployment of CEI. Further, we would like to highlight that the obtained bounds under both settings are not direct extensions from current unconstrained results. We explain the technical challenges and how we address them in Section 3. Our contributions are summarized as follows:
  • Under the frequentist setting, we derive simple regret upper bounds of $\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)\right)$ for the squared exponential (SE) kernel and $\mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)\right)$ for Matérn kernels. These bounds are improved upon the direct extension of [16] to the constrained case (see Theorem 3.8).
  • Under the Bayesian setting, we achieve the same simple regret upper bounds with high probabilities. These bounds are established based on the newly derived bounds (see Theorem 3.13) on the difference between the improvement function and its corresponding EI in the Bayesian setting.
@@ 2 Background
CBO consists of two components: the GP surrogates for the black-box objective function $f$ and constraint function $c$ and the constrained acquisition function as the sequential sampling rule guiding for the global optimum.
@@ 2.1 Gaussian process models for $f$ and $c$
Without losing generality, let the mean function for the objective GP model prior be 00 and the covariance function (kernel) be $k_{f}({\bm{x}},{\bm{x}}^{\prime}):\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$. At sample point ${\bm{x}}_{t}\in C$, we denote the objective function value as $f({\bm{x}}_{t})$ and the observed constraint function value is $c({\bm{x}}_{t})$. Given $t$ sample points, denote ${\bm{x}}_{1:t}=[{\bm{x}}_{1},\dots,{\bm{x}}_{t}]$ and ${\bm{f}}_{1:t}=[f({\bm{x}}_{1}),\dots,f({\bm{x}}_{t})]$. Moreover, denote the $t\times t$ covariance matrix ${\bm{K}}_{t}^{f}=[k_{f}({\bm{x}}_{1},{\bm{x}}_{1}),\dots,k_{f}({\bm{x}}_{1},{\bm{x}}_{t});\dots;k_{f}({\bm{x}}_{t},{\bm{x}}_{1}),\dots,k_{f}({\bm{x}}_{t},{\bm{x}}_{t})]$. The posterior distribution of $f({\bm{x}})|{\bm{x}}_{1:t},{\bm{f}}_{1:t}\sim\mathcal{N}(\mu_{t}^{f}({\bm{x}}),(\sigma_{t}^{f}({\bm{x}}))^{2})$ can then be inferred using Bayes’ rule as follows
$\mu_{t}^{f}({\bm{x}})\ =\ {\bm{k}}_{t}^{f}({\bm{x}})({\bm{K}}_{t}^{f})^{-1}{\bm{f}}_{1:t},$ (2)
$(\sigma_{t}^{f})^{2}({\bm{x}})\ =\ k^{f}({\bm{x}},{\bm{x}})-({\bm{k}}_{t}^{f}({\bm{x}}))^{T}({\bm{K}}_{t}^{f})^{-1}{\bm{k}}_{t}^{f}({\bm{x}})\ ,$
@@
$\mu_{t}^{c}({\bm{x}})\ =\ {\bm{k}}_{t}^{c}({\bm{x}})({\bm{K}}_{t}^{c})^{-1}{\bm{c}}_{1:t},$
$(\sigma_{t}^{c})^{2}({\bm{x}})\ =\ k^{c}({\bm{x}},{\bm{x}})-({\bm{k}}_{t}^{c}({\bm{x}}))^{T}({\bm{K}}_{t}^{c})^{-1}{\bm{k}}_{t}^{c}({\bm{x}})\ ,$
where ${\bm{k}}_{t}^{c}({\bm{x}})=[k_{c}({\bm{x}}_{1},{\bm{x}}),\dots,k_{c}({\bm{x}}_{t},{\bm{x}})]^{T}$, and $\mu^{c}_{t}({\bm{x}})$ and $(\sigma^{c}_{t})^{2}({\bm{x}})$ are the posterior mean and variance for $c$, respectively. Here we use the superscripts of f and c to distinguish between the GPs for $f$ and $c$.
Choices of the kernels $k_{f}$ and $k_{c}$ include the SE function, Matérn functions, etc. SE and Matérn kernels are among the most popular kernels for GP and BO. Their definitions are as follows.
@@ 2.2 Constrained Expected Improvement
Acquisition functions are critical to the performances of BO algorithms. In the unconstrained setting, one of the most widely adopted acquisition functions is EI [2]. Given $t$ samples, the improvement function of $f$ used in EI is defined as
@@
where $f^{+}_{t}=\underset{\begin{subarray}{c}i=1,\dots,t\end{subarray}}{\text{min}}\ f({\bm{x}}_{i})$. The expectation of (3) conditioned on existing samples is EI, which has a closed form [19]:
@@
Taking into account the constraint, the constrained improvement function in CEI [1] is defined as
@@
where $\Delta^{c}_{t}\in\{0,1\}$ is the feasibility indicator function where $\Delta^{c}_{t}({\bm{x}})=1$ if $c({\bm{x}})\leq 0$ and $\Delta^{c}_{t}({\bm{x}})=0$ otherwise. The incumbent $f_{t}^{+}$ in CEI is augmented to be the best feasible observation. CEI assumes that $f$ and $c$ are conditionally independent [1]. Taking the conditional expectation of (6), the CEI function is
@@
6:     Update the GP model with the addition of ${\bm{x}}_{t+1}$, $f({\bm{x}}_{t+1})$, and $c({\bm{x}}_{t+1})$.
@@
CEI can be extended to multiple constraints assuming conditional independence among the constraints [1]. Our derived convergence rates can also be readily extended to multiple constraints, as we explain in Remark 3.15.
@@ 3 Convergence rates of CEI
We present our main results of convergence rates for CEI by establishing the simple regret upper bounds. Denote the optimal solution to the constrained optimization problem (1) as ${\bm{x}}^{*}$. In the unconstrained case, the simple regret of EI is defined as $f^{+}_{t}-f({\bm{x}}^{*})$ [16]. In the constrained case, we require the current best observation to be feasible when compared to the optimal solution $f({\bm{x}}^{*})$, since one could have an infeasible sample point with smaller objecitve than $f({\bm{x}}^{*})$. Given that $f^{+}_{t}$ is already defined as the best feasible observation till iteration $t$ in CEI, we continue to use
@@
as the simple regret for CEI.
Remark 3.1 (Infeasible initial samples).
It is well-known that one potential issue with CEI is the requirement of the initial feasible sample [1]. That is, $f_{t}^{+}$ exists from the initial samples so that the CEI calculation can proceed. Recognizing this issue, solutions have been proposed including effective heuristics when $f_{t}^{+}$ does not exist [20] and using the constraint violation function to first find a feasible sample [21]. In our analysis, we make the same underlying assumption as CEI that $f_{t}^{+}$ exists.
We first establish the convergence rate under the frequentist assumptions in Section 3.1, including an improved version of the rate under frequentist assumptions in Section 3.1.1. Then, we establish the convergence rate under Bayesian assumptions in Section 3.2.
@@ 3.1 Simple regret upper bound under frequentist assumptions
In this section, we present the simple regret upper bound for CEI under the frequentist setting. Moreover, by adopting the information theory-based bounds and techniques in the noise-free cumulative regret bound of upper confidence bound (UCB) [22], we derive an improved upper bound compared to [16]. The definition of RKHS is given below.
@@
The main challenge in establishing a simple regret upper bound for CEI is how to incorporate the constraint $c({\bm{x}})\leq 0$ and the probability of feasibility function $P_{t}({\bm{x}})$ into the analysis. Existing regret bounds analysis on CBO methods often focus on UCB-type methods [23, 24], for which the acquisition functions do not have the multiplicative structure between the objective and the constraint, as we pointed out in Section 2.2.
Under Assumption 3.3, both $f$ and $c$ are bounded on $C$ by their RKHS norm bound, as stated in Lemma B.1. The simple regret upper bound is given in the following theorem.
@@
$r_{t}\leq c_{\tau B}\frac{1}{\Phi(-B_{c})}\left[B_{f}\frac{4}{t-2}+(0.4+B_{f})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (10)
@@
We start by noticing that the sum of the difference between consecutive best feasible observations is bounded, i.e., $\sum_{t=1}^{T}f^{+}_{t-1}-f^{+}_{t}\leq 2B_{f}$. Then, we adopt a technique in [16] to find $t_{k}$ such that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$, where $k\leq t_{k}\leq 2k$ and $2k\leq t\leq 2(k+1)$. Next, using the monotonicity of $f^{+}_{t}$, $r_{t}$ is bounded by $r_{t_{k}}$. Using the inequality between $I_{t}^{f}$ and $EI_{t}^{f}$ in Lemma B.4, we can bound $r_{t_{k}}$ by the EI on objective: $EI_{t_{k}}^{f}({\bm{x}}^{*})$. Then, we transform $EI_{t_{k}}^{f}({\bm{x}}^{*})$ into $EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ by inserting the term $P_{t_{k}}({\bm{x}}^{*})$, taking advantage of the multiplicative structure of CEI. The upper bound of $r_{t}$ then consists of the term $\frac{1}{P_{t_{k}}({\bm{x}}^{*})}EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. From the bound on $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$ (Lemma B.2) and the fact that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$, we can bound $EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. The constraint term $\frac{1}{P_{t_{k}}({\bm{x}}^{*})}$ remains to be bounded. We use the bound on $|c({\bm{x}})-\mu_{t}^{c}({\bm{x}})|$ in Lemma B.2 at ${\bm{x}}^{*}$ and the fact that ${\bm{x}}^{*}$ is a feasible solution to obtain a lower bound for $P_{t_{k}}({\bm{x}}^{*})$. This concludes the proof.
@@
The terms derived from the constraint function in (10) is $\frac{1}{\Phi(-B_{c})}$, which emerges from the probability of feasibility function and the posterior mean $\mu_{t}^{c}({\bm{x}})$ and standard deviation $\sigma_{t}^{c}({\bm{x}})$ of the GP model of $c({\bm{x}})$. Thanks to the multiplicative structure between the objective and constraint in $I_{t}^{C}$ (6) and $EI_{t}^{c}$ (7), the simple regret upper bound maintains a similar form.
It is clear from (10) that the convergence of $r_{t}$ relies on the posterior standard deviation $\sigma_{t}^{f}({\bm{x}}_{t+1})$. Since $t_{k}$ increases with $t$, as $\sigma_{t}^{f}({\bm{x}}_{t+1})\to 0$, so does $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. In the noise-free setting, the posterior variance can be bounded via the maximum distance between sample points and a given point. To obtain the rate of simple regret bound, we use Assumptions (1)-(4) in [16] and focus on squared exponential (SE) and Matérn kernels. Recall that the smoothness parameter of the Matérn kernel is $\nu>0$. Both the SE and Matérn kernels satisfy Assumptions (1)-(4) in [16], with SE kernel obtained as $\nu\to\infty$. Further, define
$\eta=\begin{cases}\alpha,\ \nu\leq 1\\ 0,\ \nu>1,\end{cases}$ (11)
@@
holds for at most $k$ times, for $\forall k\in\mathbb{N}$, $k\leq t$, and $i=1,\dots,t-1$. For Matérn kernels,
@@
In the constrained setting, we are able to obtain the same rates (14) as those in the unconstrained case [16] using Lemma 3.6.
Theorem 3.7.
@@
Theorem 3.7 shows that the CEI algorithm is guaranteed to find the best feasible point asymptotically with the rates elaborated in (14). Also, we point out that the choice of kernels and their parameters affect the convergence rates. Since the SE kernel can be viewed as a Matérn kernel with $\nu\to\infty$, its convergence rate is better than Matérn kernels with $\nu\leq 1$. However, due to the limitations of the kernel analysis in [16] (see Remark 3.9), for $\nu\geq 1$, SE and Matérn kernels have similar convergence rates in Theorem 3.7. As we present in the following section, improved rates for both kernels can be obtained.

### v2

@@ Abstract
Constrained Bayesian optimization (CBO) methods have seen significant success in black-box optimization with constraints. One of the most commonly used CBO methods is the constrained expected improvement (CEI) algorithm. CEI is a natural extension of expected improvement (EI) when constraints are incorporated. However, the theoretical convergence rate of CEI has not been established. In this work, we study the convergence rate of CEI by analyzing its simple regret upper bound. First, we show that when the objective function $f$ and constraint function $c$ are assumed to each lie in a reproducing kernel Hilbert space (RKHS), CEI achieves the convergence rates of $\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)\right)\ \text{and }\ \mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)\right)$ for the commonly used squared exponential and Matérn kernels ($\nu>\frac{1}{2}$), respectively. Second, we show that when $f$ is assumed to be sampled from Gaussian processes (GPs), CEI achieves similar convergence rates with a high probability. Numerical experiments are performed to validate the theoretical analysis.
@@ 1 Introduction
Bayesian optimization (BO) is an efficient method for optimizing expensive black-box functions without derivatives. It leverages probabilistic surrogate models, most commonly Gaussian processes (GPs), to balance exploration and exploitation in the search for optimal solutions (Frazier, 2018). BO has found widespread success in diverse fields such as structural design (Mathern et al., 2021), machine learning hyperparameter tuning (Wu et al., 2019), additive manufacturing process design (Wang and Papadopoulos, 2025), fusion design (Wang et al., 2024), etc.
@@
While traditional BO is typically applied to unconstrained settings, many real-world problems involve black-box constraints that must be satisfied. This has motivated growing interest in constrained Bayesian optimization (CBO), where surrogate models are also constructed for constraint functions (Bernardo et al., 2011) that are complex and expensive to evaluate, making CBO especially valuable in applications like engineering design (Song et al., 2024) and automated machine learning  (Ungredda and Branke, 2024). One of the very key difference between unconstrained and constrained optimization is that the feasible region for constrained optimization problem consists of the search space where all constraints must be satisfied. A general form of the constrained BO problem is:
$\underset{{\bm{x}}\in C}{\text{minimize}}$ $f({\bm{x}}),\$subject to $c({\bm{x}})\leq 0,$ (1)
@@
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is the objective function, and $c:\mathbb{R}^{d}\to\mathbb{R}^{m}$ are the constraint functions. Both are defined on a compact input space $C\subset\mathbb{R}^{d}$. The objective and the constraint functions are both expensive black-box functions, that can only be evaluated through expensive physical or computer experiments. Throughout this paper, we consider the noise-free setting for both the objective and the constraints, i.e., the function evaluations are deterministic and the true function values can be observed (see Remark 3.14 for discussion on the noisy case). In addition, a single constraint is considered, i.e., $m=1$, for simplicity of presentation. We note that our analysis can be easily extended to multiple constraints (see Remark 3.13 for details).
@@
Broadly, CBO methods can be categorized into implicit and explicit approaches (Amini et al., 2025). Implicit methods modify standard acquisition functions to incorporate constraints via merit functions or feasibility weights. Explicit methods estimate the feasible region directly and restrict search to this region. Among these, the constrained expected improvement (CEI) (Schonlau et al., 1998; Gelbart et al., 2014; Gardner et al., 2014) stands out as one of the most basic and widely adopted methods. CEI is a natural extension of the well-known expected improvement (EI) function (Jones et al., 1998), where the acquisition function is computed as the product of EI and the probability of feasibility. Thanks to this simple and interpretable formulation, CEI has been successfully applied across domains, and it remains one of the default choices in many constrained BO software packages  (Balandat et al., 2020).
@@
Despite its empirical popularity, the theoretical understanding of CEI lags behind. In contrast, unconstrained EI has been more extensively studied. Under a frequentist assumption where the objective $f$ lies in a reproducing kernel Hilbert space (RKHS), Bull (2011) established the convergence rate of EI by deriving the simple regret upper bound. Other works explored the density of sampled sequences (Vazquez and Bect, 2010) or connections between EI and optimal computing budget allocation (Ryzhov, 2016). However, convergence rates (i.e., simple regret upper bound) for CEI have not been rigorously established—neither under frequentist nor under Bayesian settings. Here, Bayesian setting means the objective $f$ is a function sampled from a GP.
@@
In this paper, we provide the first theoretical convergence rates for CEI, focusing on simple regret upper bounds under both the frequentist and Bayesian settings. Our convergence rates provide practitioners theoretical assurance for the practical deployment of CEI. We explain the technical challenges and how we address them in Section 3. Our contributions are summarized as follows:
  • Under the frequentist setting, we derive simple regret upper bounds of $\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)\right)$ for the squared exponential (SE) kernel and $\mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)\right)$ for Matérn kernels ($\nu>\frac{1}{2}$). These bounds are improved upon the direct extension of Bull (2011) to the constrained case for SE kernel with $d\geq 3$ and Matérn kernels with $d\geq 3,\nu\geq\frac{d}{d-2}$. (see Theorem 3.7).
  • Under the Bayesian setting for the objective, we achieve similar simple regret upper bounds with high probabilities. These bounds are established based on the newly derived bounds (see Theorem 3.11) on the difference between the improvement function and its corresponding EI in the Bayesian setting.
@@ 2 Background
CBO mainly consists of two components: the GP surrogates for the black-box objective function $f$ and constraint function $c$, and the constrained acquisition function as the sequential sampling rule guiding for the global optimum.
@@ 2.1 Gaussian process models for $f$ and $c$
Without losing generality, let the mean function for the objective GP model prior be 0 and the covariance function (kernel) be $k_{f}({\bm{x}},{\bm{x}}^{\prime}):\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$. At sample point ${\bm{x}}_{t}\in C$, we denote the objective function value as $f({\bm{x}}_{t})$ and the observed constraint function value is $c({\bm{x}}_{t})$. Given $t$ sample points, denote ${\bm{x}}_{1:t}=[{\bm{x}}_{1},\dots,{\bm{x}}_{t}]$ and ${\bm{f}}_{1:t}=[f({\bm{x}}_{1}),\dots,f({\bm{x}}_{t})]$. Moreover, denote the $t\times t$ covariance matrix ${\bm{K}}_{t}^{f}=[k_{f}({\bm{x}}_{1},{\bm{x}}_{1}),\dots,k_{f}({\bm{x}}_{1},{\bm{x}}_{t});\dots;k_{f}({\bm{x}}_{t},{\bm{x}}_{1}),\dots,k_{f}({\bm{x}}_{t},{\bm{x}}_{t})]$. The posterior distribution of $f({\bm{x}})|{\bm{x}}_{1:t},{\bm{f}}_{1:t}\sim\mathcal{N}(\mu_{t}^{f}({\bm{x}}),(\sigma_{t}^{f}({\bm{x}}))^{2})$ can then be inferred using Bayes’ rule as follows
$\mu_{t}^{f}({\bm{x}})\ =\ ({\bm{k}}_{t}^{f}({\bm{x}}))^{T}({\bm{K}}_{t}^{f})^{-1}{\bm{f}}_{1:t},$ (2)
$(\sigma_{t}^{f})^{2}({\bm{x}})\ =\ k_{f}({\bm{x}},{\bm{x}})-({\bm{k}}_{t}^{f}({\bm{x}}))^{T}({\bm{K}}_{t}^{f})^{-1}{\bm{k}}_{t}^{f}({\bm{x}})\ ,$
@@
$\mu_{t}^{c}({\bm{x}})\ =\ ({\bm{k}}_{t}^{c}({\bm{x}}))^{T}({\bm{K}}_{t}^{c})^{-1}{\bm{c}}_{1:t},$
$(\sigma_{t}^{c})^{2}({\bm{x}})\ =\ k_{c}({\bm{x}},{\bm{x}})-({\bm{k}}_{t}^{c}({\bm{x}}))^{T}({\bm{K}}_{t}^{c})^{-1}{\bm{k}}_{t}^{c}({\bm{x}})\ ,$
where ${\bm{k}}_{t}^{c}({\bm{x}})=[k_{c}({\bm{x}}_{1},{\bm{x}}),\dots,k_{c}({\bm{x}}_{t},{\bm{x}})]^{T}$, and $\mu^{c}_{t}({\bm{x}})$ and $(\sigma^{c}_{t})^{2}({\bm{x}})$ are the posterior mean and variance for $c$, respectively. Here we use the subscripts f, c and superscripts f, c to distinguish between GPs for $f$ and $c$. Choices of the kernels $k_{f}$ and $k_{c}$ include the SE and Matérn kernels, which are among the most popular kernels for GP and BO. Their definitions are as follows.
@@ 2.2 Constrained Expected Improvement
Acquisition functions are critical to the performances of BO algorithms. In the unconstrained setting, one of the most widely adopted acquisition functions is EI (Jones et al., 1998). Given $t$ samples, the improvement function of $f$ used in EI is defined as
@@
where $f^{+}_{t}=\underset{\begin{subarray}{c}i=1,\dots,t\end{subarray}}{\text{min}}\ f({\bm{x}}_{i})$. The expectation of (3) conditioned on existing samples is EI, which has a closed form Brochu et al. (2010):
@@
Taking into account the constraint, the constrained improvement function in CEI (Gardner et al., 2014) is defined as
@@
where $\Delta^{c}_{t}\in\{0,1\}$ is the feasibility indicator function where $\Delta^{c}_{t}({\bm{x}})=1$ if $c({\bm{x}})\leq 0$ and $\Delta^{c}_{t}({\bm{x}})=0$ otherwise. The incumbent $f_{t}^{+}$ in CEI is augmented to be the best feasible observation. CEI assumes that $f$ and $c$ are conditionally independent (Gardner et al., 2014). Taking the conditional expectation of (6), the CEI function is
@@
6:  Update the GP models with the addition of ${\bm{x}}_{t+1}$, $f({\bm{x}}_{t+1})$, and $c({\bm{x}}_{t+1})$.
@@
CEI can be extended to multiple constraints assuming conditional independence among the constraints (Gardner et al., 2014). Our derived convergence rates can also be readily extended to multiple constraints, as we explain in Remark 3.13.
@@ 3 Convergence rates of CEI
We present our main results of convergence rates for CEI by establishing the simple regret upper bounds. Denote the optimal solution to the constrained optimization problem (1) as $\bm{x}$$^{*}$. In the unconstrained case, the simple regret of EI is defined as $f^{+}_{t}-f({\bm{x}}^{*})$ (Bull, 2011). In the constrained case, we use the current best feasible observation and compare it to the optimal solution $f({\bm{x}}^{*})$, since one could have an infeasible sample point with smaller objective than $f({\bm{x}}^{*})$. Given that $f^{+}_{t}$ is already defined as the best feasible observation till iteration $t$ in CEI, we continue to use
@@
as the simple regret for CEI. In our analysis, we make the same underlying assumption as CEI that $f_{t}^{+}$ exists. In the following, we first establish the convergence rate under the frequentist assumptions in Section 3.1, including an improved version of the rate under frequentist assumptions in Section 3.1.1. Then, we establish the convergence rate under Bayesian objective assumptions in Section 3.2.
@@ 3.1 Simple regret upper bound under frequentist assumptions
In this section, we present the simple regret upper bound for CEI under the frequentist setting. Moreover, by adopting the information theory-based bounds and techniques in the noise-free cumulative regret bound of upper confidence bound (UCB) (Lyu et al., 2019), we can derive an improved upper bound in some cases compared to Bull (2011). The definition of RKHS is given below.
@@
The main challenge in establishing a simple regret upper bound for CEI is how to incorporate the constraint $c({\bm{x}})\leq 0$ and the probability of feasibility function $P_{t}({\bm{x}})$ into the analysis. Existing regret bounds analysis on CBO methods often focus on UCB-type methods (Lu and Paulson, 2022; Zhou and Ji, 2022), for which the acquisition functions do not have the multiplicative structure between the objective and the constraint.
Under Assumption 3.2, both $f$ and $c$ are bounded on $C$ by their RKHS norm bounds, as stated in Lemma B.1. The simple regret upper bound is given in the following theorem.
@@
$r_{t}\leq\frac{c_{\tau B}}{\Phi(-B_{c})}\left[B_{f}\frac{4}{t-2}+(0.4+B_{f})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (10)
@@
We start by noticing that the sum of the difference between consecutive best feasible observations is bounded, i.e., $\sum_{t=1}^{T}f^{+}_{t-1}-f^{+}_{t}\leq 2B_{f}$. Then, we adopt a technique in Bull (2011) to find $t_{k}$ such that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$, where $k\leq t_{k}\leq 2k$ and $2k\leq t\leq 2(k+1)$. Next, using the monotonicity of $f^{+}_{t}$, $r_{t}$ is bounded by $r_{t_{k}}$. Using the inequality between $I_{t}^{f}$ and $EI_{t}^{f}$ in Lemma B.4, we can bound $r_{t_{k}}$ by the EI on objective: $EI_{t_{k}}^{f}({\bm{x}}^{*})$. Then, we transform $EI_{t_{k}}^{f}({\bm{x}}^{*})$ into $EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ by inserting the term $P_{t_{k}}({\bm{x}}^{*})$, taking advantage of the multiplicative structure of CEI. The upper bound of $r_{t}$ then consists of the term $\frac{1}{P_{t_{k}}({\bm{x}}^{*})}EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. From the confidence interval $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$ (Lemma B.2) and the fact that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$, we can bound $EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. The constraint term $\frac{1}{P_{t_{k}}({\bm{x}}^{*})}$ remains to be bounded. We use the confidence interval on $|c({\bm{x}})-\mu_{t}^{c}({\bm{x}})|$ in Lemma B.2 at $\bm{x}$$^{*}$ and the fact that $\bm{x}$$^{*}$ is a feasible solution to obtain a lower bound for $P_{t_{k}}({\bm{x}}^{*})$. This concludes the proof.
@@
The terms derived from the constraint function in (10) is $\frac{1}{\Phi(-B_{c})}$, which emerges from the probability of feasibility function and $\mu_{t}^{c}({\bm{x}})$ and $\sigma_{t}^{c}({\bm{x}})$ of the GP model of $c({\bm{x}})$. Thanks to the multiplicative structure between the objective and constraint in $I_{t}^{C}$ (6) and $EI_{t}^{c}$ (7), the simple regret upper bound maintains a similar form.
It is clear from (10) that the convergence of $r_{t}$ relies on the posterior standard deviation $\sigma_{t}^{f}({\bm{x}}_{t+1})$. Since $t_{k}$ increases with $t$, as $\sigma_{t}^{f}({\bm{x}}_{t+1})\to 0$, so does $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. In the noise-free setting, the posterior variance can be bounded via the maximum distance between sample points and a given point. To obtain the rate of simple regret bound, we use Assumptions (1)-(4) in Bull (2011) and focus on squared exponential (SE) and Matérn kernels. Recall that the smoothness parameter of the Matérn kernel is $\nu>0$. Both the SE and Matérn kernels satisfy Assumptions (1)-(4) in Bull (2011), with SE kernel obtained as $\nu\to\infty$. Further, define
[…]
@@
holds for at most $k$ times, for $\forall k\in\mathbb{N}$, $k\leq t$ and $i=1,\dots,t-1$. For Matérn kernels,
@@
In the constrained setting, we are able to obtain the same rates as those in the unconstrained case (Bull, 2011) using Lemma 3.5.
Corollary 3.6.
@@
Corollary 3.6 shows that the CEI algorithm is guaranteed to find the best feasible point asymptotically with the rates elaborated in (14). Also, we point out that the choice of kernels and their parameters affect the convergence rates. Since the SE kernel can be viewed as a Matérn kernel with $\nu\to\infty$, its convergence rate is better than Matérn kernels with $\nu\leq 1$. However, due to the limitations of the kernel analysis in Bull (2011) (see Remark 3.8), for $\nu\geq 1$, SE and Matérn kernels have similar convergence rates in Corollary 3.6. As we present in the following section, improved rates for both kernels can be obtained in some cases.